what is the answers to:
13-4y=25
3/4x-5>7
-1/2x-7=23

Answer by rfer(16322)

13-4y=25
-4y=12
y=-3
------------
3/4x-5 > 7
3/4x > 7+5
3/4x > 13
x > 13*4/3
x > 52/3
x > 17 1/3
---------------
-1/2x-7=23
-1/2x=30
x=30*-2/1
x=-60